Question 24

Consider the statement 

If n is a multiple of 3, then 2n + 2 is not a multiple of 3.

The converse of the given statement is:

 If n is not a multiple of 3, then 2n + 2 is a multiple of 3.

1. True

2. False

Question 25

Consider the following statement, for all x  Z:

 If x + 1 is even, then 3x2

- 4 is odd. 

The correct way to start a direct proof to determine if the statement is true is as follows:

 Assume x is even, then x = 2k for some k  Z,

 then 3x2 – 4

 ie 3(2k)2

- 4

 ie ………..

1. True

2. False


1
Expert's answer
2021-07-26T06:42:46-0400

Question 24


Consider the statement "If "n" is a multiple of 3, then "2n + 2" is not a multiple of 3". The converse of a statement is formed by switching the hypothesis and the conclusion. Hence, the converse of the given statement is "If "2+2n" is not a multiple of 3, then "n" is a multiple of 3".


Answer: 2. False



Question 25


Consider the following statement, "for all "x \in\Z" : If "x + 1" is even, then "3x^2- 4" is odd. 

The correct way to start a direct proof to determine if the statement is true is as follows:

 Assume "x+1" is even, then "x +1= 2k" for some "k \in \Z",  then "3x^2 \u2013 4=3(2k-1)^2 \u2013 4" ...


Answer: 2. False